# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v1

@@ 4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (et al., 2019) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (Panayotov et al., 2015) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
@@ 4.1.1. ASVSpoof 2019
The ASVSpoof 2019 dataset is the primary benchmark for evaluating spoofing countermeasures. We use the Logical Access (LA) portion, whose partitions are statistically summarized in Table 3. To create a robust and challenging test set for our experiments against anti-spoofing systems, we curated a specific subset from the official Evaluation partition. As shown in the table, we selected 100 voice samples for each of the 13 advanced spoofing attack types (A07–A19), resulting in a balanced test set of 1,300 spoofed samples for our analysis.
@@
Partition Speakers (M/F) Total Speakers Total Utt. Selected (per Attack) Total Selected
@@ 4.1.2. LibriSpeech
For our end-to-end experiments evaluating the full pipeline, we used the dev-clean and test-clean subsets of the LibriSpeech ASR corpus. As detailed in Table 4, these subsets provide high-quality audio from a total of 80 unique speakers, evenly distributed between males and females. To simulate a realistic user verification scenario, we selected a specific number of samples for each speaker. From each of the 80 speakers, we designated 5 utterances for enrollment to create their voiceprint and a separate set of 5 utterances for verification trials. This resulted in a total of 400 selected audio files from each subset.
Table 4. Statistics of the LibriSpeech Subsets and Experimental Selection.
Subset Speakers (M/F) Total Speakers Utterances Selected (per User) Total Selected
dev-clean 20 / 20 40 2,703 10 (5 enroll, 5 verify) 400
test-clean 20 / 20 40 2,620 10 (5 enroll, 5 verify) 400
@@ 4.2. Authentication Models
To ensure a comprehensive evaluation, we test our attack against a diverse set of three state-of-the-art (SOTA) Speaker Verification (SV) systems. This set includes two well-known open-source models and one commercial black-box API, allowing us to assess performance on both accessible academic benchmarks and proprietary, real-world systems.
  • X-Vectors (Snyder et al., 2018): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (Ravanelli et al., 2021) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (Li et al., 2017)  [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (Microsoft, [n. d.])  [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
@@ 4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (et al., 2019). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (Tak et al., 2021b)  [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (et al., 2019). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (Tak et al., 2021a)  [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (et al., 2019).
  • RawPC-DARTS (Ge et al., 2021)  [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (et al., 2019), making it a critical system to evaluate against.
@@ 5. Evaluation Methodology
5.1. End-to-End System Evaluation on ASVspoof 2019
To ensure our methodology is grounded in a standard framework, we conduct our experiments on the ASVspoof 2019 dataset (et al., 2019). This dataset is a well-established and widely recognized benchmark for research on voice spoofing and countermeasures. While we do not perform a direct comparison with existing literature in this study, our use of this dataset provides a crucial baseline. This approach allows future work to easily and effectively utilize our results for a direct comparison with our attack. In this setup, we use the genuine bona fide audio samples for speaker enrollment and the dataset’s pre-existing spoofed audio files as the base for our perturbations.
5.2. End-to-End System Evaluation on LibriSpeech
Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (Panayotov et al., 2015). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (Liao et al., 2024). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
5.3. Targeted System Evaluations
To gain deeper insights into the vulnerabilities of individual system components, we conduct targeted experiments against CMs and VAS separately.
  • Attacks on Anti-Spoofing Systems using LibriSpeech: We first perform a dedicated experiment targeting only the CMs. As described above, these test audios were generated from a pair combination of different prompts and LibriSpeech audios (Panayotov et al., 2015) that passed to Fish Speech APIs (Liao et al., 2024), then passed to our SMIA attack to generate adversarial audios that tested against different CMs to test their vulnerability.
  • Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019: To ensure a direct and fair comparison with prior work, we conduct a second experiment targeting CMs using the ASVspoof 2019 dataset (et al., 2019). We used the spoofed audios from the ASVspoof 2019 dataset (et al., 2019), then passed them to our SMIA framework to test them against different CMS. This methodology allows for a clear performance benchmark against the results reported by both Kassis et al. (Kassis and Hengartner, 2023) and Hai et al. (Hai et al., 2023).
  • Attacks on Voice Authentication Systems: We conduct an extensive evaluation to demonstrate the significant vulnerability of VAS. Using the LibriSpeech (Panayotov et al., 2015) and Fish Speech (Liao et al., 2024), we test our generated audio against the VAS models across three distinct biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
  • Computational Efficiency Evaluation: To provide a complete picture, we evaluated SMIA’s computational efficiency against SiFDetectCracker (Hai et al., 2023). We measured both the number of iterations and the total time needed to carry out a successful attack. This analysis confirms our framework’s practical feasibility, highlighting its significant speed advantage over prior work.
  • Robustness in Real-World Scenarios: An effective attack must be robust to the distortions introduced by real-world transmission channels. To evaluate our method’s practicality, we test the generated audio against the VAS systems in two simulated environments: over-the-air, which mimics audio being played through a speaker and recorded by a microphone, and over-the-line, which simulates the effects of a VoIP codec. This demonstrates the attack’s applicability beyond controlled digital environments.
5.4. Stealth and Detectability Analysis
To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative analysis. We show that while adversarial audio generated by methods, like SiFDetectCracker (Hai et al., 2023) can be easily identified and neutralized by a simple high-pass filter, our attack audio remains effective. This highlights our method’s superior stealth and its ability to bypass rudimentary digital forensic defenses.
5.5. Ablation Study
To understand which parts of our attack contribute most to its success, we conducted two specific ablation studies.
@@ 6. Results
This section presents the comprehensive evaluation of our proposed Spectral Masking & Interpolation Attack (SMIA). We analyze its effectiveness against state-of-the-art anti-spoofing countermeasures (CMs) and voice authentication systems (VAS) across the various scenarios outlined in our methodology.
6.1. End-to-End System Evaluation on ASVspoof 2019
The results of our evaluation on the ASVspoof 2019 dataset, a standardized benchmark for voice security research, are presented in Table 5. Crucially, our findings demonstrate that the SMIA attack is highly effective even against these layered defenses, achieving a formidable ASR of up to 99.3% against combined VAS and CM security pipelines.
A direct numerical comparison with prior works, such as Kassis et al. (Kassis and Hengartner, 2023), was not conducted, as they often report different metrics that are not directly comparable to the ASR used in our analysis. Similarly, a comparison with Hai et al. (Hai et al., 2023) is not included, as their analysis did not extend to the combined VAS and CM systems that are the focus of our evaluation. Therefore, the purpose of this experiment is not to draw a direct comparison, but to establish a robust and reproducible performance baseline for the SMIA attack. These results provide a clear benchmark for any future research conducted on this standard dataset.
Table 5. Attack Success Rate (ASR) (%) of SMIA against end-to-end systems (VAS and CMs) on the ASVspoof 2019 dataset.
VAS Model Anti-Spoofing CM
@@ 6.1. End-to-End System Evaluation on ASVspoof 2019
X-Vectors 99.3 80 99.3
DeepSpeaker 67.9 47.3 69.6
@@ 6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (Panayotov et al., 2015), Table 6 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations and never falling below a formidable 82.7% success rate. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses and underscoring the urgent need for more resilient countermeasures.
@@ 6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (et al., 2019). The results, presented in Table 7, are unequivocal. SMIA substantially outperforms both state-of-the-art baselines across all tested countermeasures, achieving a near-perfect 99.3% ASR and demonstrating a significant advancement in attack efficacy.
Table 7. Performance comparison (ASR) of SMIA against state-of-the-art attacks on the ASVspoof 2019 dataset.
@@
SMIA 99.3 81.1 99.3
SiFDetectCracker (Hai et al., 2023) 80.4 75.8 84.1
Kassis et al. (Kassis and Hengartner, 2023) 12.03 15.82
@@ 6.3.2. Attacks on Anti-Spoofing Systems using LibriSpeech
A subsequent experiment using adversarial audio from the LibriSpeech pipeline revealed catastrophic vulnerabilities across all tested CMs. As detailed in Table 8, the results were definitive: our SMIA attack achieved a perfect 100% ASR against every state-of-the-art model evaluated.
@@ 6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 9, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (Li et al., 2017), X-Vectors (Snyder et al., 2018), and the commercial Microsoft Azure SV (Microsoft, [n. d.]), peaking at a perfect 100%.
@@
Attack Task Model Recall ASR (%) F1-score EER
@@
Kassis et al. (Kassis and Hengartner, 2023) SV X-Vectors - 100 - -
@@ 6.3.4. Computational Efficiency
For a holistic evaluation, we further compared the time cost of SMIA against the baselines. As summarized in Table 10, our framework consistently required fewer iterations and lower computational time to achieve successful attacks. Specifically, SMIA achieved near-perfect attack success within minutes, while Hai et al. (Hai et al., 2023) required significantly longer runtimes under comparable settings. This reduction in time cost emphasizes the practical efficiency of SMIA, making it not only more effective but also more feasible for real-world adversarial scenarios.
Table 10. Average runtime comparison of SMIA and Hai et al. (Hai et al., 2023).
Attack Avg Time (sec)
SMIA 28
Hai et al. (Hai et al., 2023) 185

### v3

@@ 4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (Todisco2019ASVspoof) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (panayotov2015librispeech) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
@@ 4.1.1. ASVSpoof 2019
The ASVSpoof 2019 dataset is the primary benchmark for evaluating spoofing countermeasures. We use the Logical Access (LA) Evaluation partition, whose partitions are statistically summarized in Table 2.
@@
Partition Speakers (M/F) Total Speakers Total Utt.
@@ 4.1.2. LibriSpeech
We used the test-clean subset of the LibriSpeech ASR corpus. This subset provides high-quality audio from 40 unique speakers (20 male and 20 female). To simulate a realistic user verification scenario, we selected a specific number of samples from this subset. From each of the 40 speakers, we designated 5 utterances for enrollment to create their voiceprint and a separate set of 5 utterances for verification trials. This resulted in a total of 400 selected audio files (10 per speaker) for our evaluation.
@@ 4.2. Authentication Models
To ensure a comprehensive evaluation, we test our attack against a diverse set of three SOTA SV systems. This set includes two well-known open-source models and one commercial black-box API, allowing us to assess performance on both accessible academic benchmarks and proprietary, real-world systems.
  • X-Vectors (xvectors): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (ravanelli2021speechbrain) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (deepspeaker) [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (microsoft_speaker) [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
@@ 4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (Todisco2019ASVspoof). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (rawnet2) [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (Todisco2019ASVspoof). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (rawgat) [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (Todisco2019ASVspoof).
  • RawPC-DARTS (rawdarts) [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (Todisco2019ASVspoof), making it a critical system to evaluate against.
@@ 5. Evaluation Methodology
    End-to-End System Evaluation on ASVspoof: To create a robust evaluation subset, we randomly selected 30 speakers from the ASVspoof 2019 LA Evaluation partition (Todisco2019ASVspoof). For each selected speaker, we used 5 of their genuine bona fide audio samples for enrollment in the VAS. We then randomly selected 5 of their corresponding pre-existing spoofed audio files to serve as the base samples for our SMIA attack, resulting in 150 attack trials (30 speakers $\times$ 5 samples).
  2. (2)
    End-to-End System Evaluation on LibriSpeech: Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (panayotov2015librispeech). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (fish-speech-v1.4). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
  3. (3)
    Targeted System Evaluations: To gain deeper insights into the vulnerabilities of individual system components, we conduct targeted experiments against CMs and VAS separately.
    1. (a)
      Benchmark Attacks on Anti-Spoofing Systems (ASVspoof): To ensure a fair comparison with prior work (kassis2023breaking; hai2023sifdetectcracker), we applied our SMIA framework to spoofed audios from the ASVspoof 2019 dataset (Todisco2019ASVspoof). This methodology allows for a direct performance benchmark against existing methods. Specifically, we targeted all 13 logical access spoofing attack types (A07-A19) present in the evaluation partition. For each attack type, we randomly selected 50 audio samples, creating a comprehensive test set of 650 spoofed utterances (13 types $\times$ 50 samples) to ensure our attack’s generalizability.
    2. (b)
      Attacks on Anti-Spoofing Systems (LibriSpeech): We conducted a dedicated experiment targeting CMs. Test audios were generated using LibriSpeech (panayotov2015librispeech) and Fish Speech APIs (fish-speech-v1.4), then processed by our SMIA attack to test the vulnerability of various CMs.
    3. (c)
      Attacks on Voice Authentication Systems (VAS): We extensively evaluated VAS vulnerabilities using audio generated from LibriSpeech (panayotov2015librispeech) and Fish Speech (fish-speech-v1.4). The attacks were tested against VAS models across three biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
    4. (d)
      Computational Efficiency Evaluation: We evaluated SMIA’s efficiency against SiFDetectCracker (hai2023sifdetectcracker), measuring the iterations and time required for a successful attack.
    5. (e)
      Robustness in Real-World Scenarios: To assess practical applicability, we tested the generated audio against VAS systems in two simulated real-world environments: over-the-air (speaker-to-microphone) and over-the-line (VoIP codec simulation).
  4. (4)
    Stealth and Detectability Analysis: To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative visual analysis of the audio spectrograms. We compare the adversarial samples generated by SMIA against those from SiFDetectCracker (hai2023sifdetectcracker) to show that our method avoids the structured, conspicuous artifacts that create a detectable pattern. This highlights our method’s superior stealth and resilience against basic forensic analysis.
  5. (5)
    Ablation Study: To understand which parts of our attack contribute most to its success, we conducted these ablation studies:
    1. (a)
@@ 6. Results
This section presents the comprehensive evaluation of our proposed SMIA. We analyze its effectiveness against SOTA CMs and VAS across the various scenarios outlined in our methodology.
6.1. End-to-End System Evaluation on ASVspoof
The results of our evaluation on the ASVspoof 2019 dataset, a standardized benchmark for voice security research, are presented in Table 3. Crucially, our findings demonstrate that the SMIA attack is highly effective even against these layered defenses, achieving a formidable ASR of up to 97% against combined VAS and CM security pipelines.
Our evaluation shows that the proposed method significantly outperforms Kassis et al. (kassis2023breaking). In contrast, a direct comparison with Hai et al. (hai2023sifdetectcracker) remains unfeasible, as their work does not address the combined VAS and CM systems utilized in our analysis. Consequently, this experiment provides evidence of superior performance where comparison is possible, while also providing a clear benchmark for future research on this specific dataset configuration.
Table 3. ASR (%) of SMIA against end-to-end systems (VAS and CMs) on the ASVspoof 2019 dataset.
Group VAS Model Anti-Spoofing CM
@@ 6.1. End-to-End System Evaluation on ASVspoof
Ours DeepSpeaker 59 71 30
X-Vectors 97 93 82
Kassis et al. X-Vectors - 14.25 9.67
@@ 6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (panayotov2015librispeech), Table 4 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses.
However, a key finding in Table 4 is the notable ASR drop to 66.5% when targeting the DeepSpeaker + RawPC-Darts combination. This is not a failure of the attack but rather a critical insight into the core tradeoff of our problem. It highlights a fundamental conflict: the perturbations required to fool the robust RawPC-Darts (CM) are strong enough to degrade the speaker’s biometric identity, causing DeepSpeaker (VAS) to fail the verification.
@@ 6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (Todisco2019ASVspoof). The results, presented in Table 5, are unequivocal. SMIA substantially outperforms both SOTA baselines across all tested countermeasures, demonstrating a significant advancement in attack efficacy.
Table 5. Performance comparison (ASR) of SMIA against SOTA attacks on the ASVspoof 2019 dataset.
@@
SMIA 96.3 95 87
SiFDetectCracker (hai2023sifdetectcracker) 80.4 75.8 84.1
Kassis et al. (kassis2023breaking) 12.03 15.82
@@ 6.3.2. Attacks on Anti-Spoofing Systems using LibriSpeech
A subsequent experiment using adversarial audio from the LibriSpeech pipeline demonstrated the effectiveness of the attack across different architectures. As detailed in Table 6, we observed a perfect 100% ASR against two of the SOTA models. For the RawDarts-PC, while not total, the attack successfully bypassed the system with an 87.9% ASR, confirming robust performance across the evaluated set.
@@ 6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 7, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (deepspeaker), X-Vectors (xvectors), and the commercial Microsoft Azure SV (microsoft_speaker), peaking at a perfect 100%.
@@
Attack Task Model Recall ASR F1-score EER
@@
Kassis et al. (kassis2023breaking) SV X-Vectors - 100 - -
@@ 6.3.4. Computational Efficiency
For a holistic evaluation, we compared the efficiency of SMIA against the SiFDetectCracker baseline (hai2023sifdetectcracker) when targeting standalone CMs. We evaluated efficiency using two key metrics: (1) the average number of model queries required for a successful attack, and (2) the average computational time. The query count is a more robust, hardware-independent metric for BBO attacks, while the time cost demonstrates practical feasibility.
As summarized in Table 8, our framework demonstrates superior computational efficiency. We achieved a slight reduction in the average query count—21.6 compared to 22.3 for the baseline. However, the most significant advantage lies in the processing speed; our attack was 6.6 times faster in terms of raw time, emphasizing its practical feasibility.
Table 8. Efficiency comparison (Avg. Queries and Time) of SMIA and SiFDetectCracker (hai2023sifdetectcracker) against standalone CMs.
Attack Avg. Queries Avg. Time (sec)
SMIA 21.6 28
SiFDetectCracker (hai2023sifdetectcracker) 22.3 185
(a) Original
(b) Hai et al. (hai2023sifdetectcracker)
(c) SMIA
(d) Original
(e) Hai et al. (hai2023sifdetectcracker)
(f) SMIA
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (hai2023sifdetectcracker) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (hai2023sifdetectcracker) (3(b), 3(e)), which create a detectable pattern. In contrast, the perturbations from SMIA (3(c), 3(f)) are more subtle and randomly distributed, making them significantly harder to detect.